### v2

# Deep learning for predicting the occurrence of tipping points

Abstract
Tipping points occur in many real-world systems, at which the system shifts suddenly from one state to another. The ability to predict the occurrence of tipping points from time series data remains an outstanding challenge and a major interest in a broad range of research fields. Particularly, the widely used methods based on bifurcation theory are neither reliable in prediction accuracy nor applicable for irregularly-sampled time series which are commonly observed from real-world systems. Here we address this challenge by developing a deep learning algorithm for predicting the occurrence of tipping points in untrained systems, by exploiting information about normal forms. Our algorithm not only outperforms traditional methods for regularly-sampled model time series but also achieves accurate predictions for irregularly-sampled model time series and empirical time series. Our ability to predict tipping points for complex systems paves the way for mitigation risks, prevention of catastrophic failures, and restoration of degraded systems, with broad applications in social science, engineering, and biology.
Introduction
Many real-world systems, ranging from biological systems to climate systems and financial systems, can experience sudden shifts between states at critical thresholds which are so-called tipping points, for example, the epileptic seizures(1), abrupt shifts in ocean circulation(2, 3) and systemic market crashes(4). Accurately predicting tipping points before they occur has many important applications, including making strategies to prevent disease outbreak, avoiding disasters induced by climate change and designing robust financial systems. However, due to the complexity of various real-world systems, it is a challenge task to invent an effective tool to predict the occurrence of tipping points(5, 6).
A classical mathematical tool to understand tipping points is bifurcation theory which focuses on how dynamical systems undergo sudden qualitative changes as a parameter crosses a threshold(7). There are basically two classes of methods based on bifurcation theory to deal with tipping points prediction. The first class of methods is using lag-1 autocorrelation which is based on “critical slowing down”, a paramount clue of whether a tipping point is approached(5, 8). Such methods are widely used in various complex systems, including degenerate fingerprinting(9), BB method(10) and ROSA(11). The second class of methods is using dynamical eigenvalue (DEV)(12) which is based on Takens embedding theorem(13). Yet, the success of the existing methods based on bifurcation theory are impaired by two fundamental limitations. First, these methods are not applicable for irregularly-sampled time series data which are commonly observed from various scientific fields, such as geoscientific measurements(14, 15), medical observations(16) and biological systems(17). Second, the performance of existing methods based on bifurcation theory in prediction accuracy is affected by several approximations used in these methods. One such approximation is that only the first-order term of the dynamical systems is considered in these methods while the impact of the higher-order terms is ignored. However, the higher-order terms become significant for prediction when a tipping point is approached(18) (Supplementary information Note S1). Besides, in the approximation of fast-slow systems(19), the delay on the bifurcation-tipping due to the changing rate of the bifurcation parameter is ignored(20, 21) (Supplementary information Note S2).
Recent studies show that some machine learning based techniques have been used as effective early warning signals of tipping points by learning generic features of bifurcation(18, 22, 23, 24). However, they can not be used for predicting where the tipping points occur. A machine learning framework based on reservoir computing has been developed for tipping points prediction(25, 26, 27). This algorithm requires time series sampled from all interacting variables of study system for training and prediction. However, this information is often not available for real-world systems. Therefore, it is often not feasible to use this machine learning framework for predicting tipping points of real-world systems.
To circumvent above limitations of existing methods for tipping points prediction, we develop a deep learning (DL) algorithm based on 2D CNN-LSTM architecture. Based on the embedding theorem for irregular sampling(28) (Supplementary information Note S3), this DL algorithm only requires the time series of the state and that of the bifurcation parameter from a single variable of a system (Fig. 1). The 2D CNN layer uses convolution kernels with length $d$ to extract features of the system reconstructed by a $d$-dimensional delay embedding. These convolutional kernels move along the time series, extracting features from each segment of the state time series and the bifurcation parameter time series, thereby generating sequences of feature representations. Then the LSTM layer is used to identify long-term dependencies from the sequences of features of the reconstructed system for tipping points prediction. We assume that the DL algorithm can detect features that emerge in time series prior to a tipping point, such as the features of the recovery rate in normal forms, which are associated with the occurrence of tipping points. We apply the DL algorithm to predict the occurrence of tipping points in systems it was not trained on. Our DL algorithm is applicable to both regularly-sampled and irregularly-sampled time series. We first test this DL algorithm on regularly-sampled time series generated by the models from ecology(29, 30, 31) and climatology(32, 33, 34, 35). The DL algorithm outperforms traditional methods in prediction accuracy for regularly-sampled model time series. We further validate our DL algorithm on irregularly-sampled time series generated by the same models and two other models from neuroscience with hysteresis phenomena(36, 37). Finally, we validate our DL algorithm in irregularly-sampled empirical data from microbiology(38) and thermoacoustics(20). In this work, we show that our DL algorithm is effective in dealing with dynamical systems exhibiting fold, Hopf and transcritical bifurcation. We anticipate that our DL algorithm may also be applicable to dynamical systems exhibiting other bifurcation types.
Fig. 1. The 2D CNN-LSTM architecture. The 2D convolutional kernel processes state series and bifurcation parameter series into one-dimensional features series of the reconstructed system $(x_{i},x_{i+1},\dots,x_{i+d-1})$. This features series is then subjected to local max pooling, followed by analysis using an LSTM layer. Finally, the features processed by the LSTM are mapped to the predicted tipping point.
Results
Bifurcation theory
The bifurcation theory is a classical and widely used mathematical tool to understand tipping points. According to the center manifold theorem(7), as a high-dimensional dynamical system approaches a bifurcation, its dynamics converges to a lower-dimensional space which exhibits dynamics topologically equivalent to those of the normal form of that bifurcation(7). Here we focused on codimension-one bifurcations in continuous-time dynamical systems, including the fold, Hopf, and transcritical bifurcation. Many tipping points of complex systems from the nature and the society are initiated by these three types of bifurcation(18, 39). The normal forms of fold, Hopf, and transcritical bifurcation are shown in Fig. 2 A-C respectively.
Suppose $dx/dt=f(x,\mu(t))$ has quasi-static attractor $x^{*}(\mu)$ where $\mu$ is the bifurcation parameter. Then the recovery rate is defined as the maximal real part of the eigenvalues of the Jacobian matrix when $x=x^{*}(\mu)$ in an $n$-dimensional dynamical system(40, 41)
$\lambda=max(Re(eigvals(\begin{bmatrix}\frac{\partial f_{1}}{\partial x_{1}}&\frac{\partial f_{1}}{\partial x_{2}}&\cdots&\frac{\partial f_{1}}{\partial x_{n}}\\ \frac{\partial f_{2}}{\partial x_{1}}&\frac{\partial f_{2}}{\partial x_{2}}&\cdots&\frac{\partial f_{2}}{\partial x_{n}}\\ \vdots&\vdots&\ddots&\vdots\\ \frac{\partial f_{n}}{\partial x_{1}}&\frac{\partial f_{n}}{\partial x_{2}}&\cdots&\frac{\partial f_{n}}{\partial x_{n}}\end{bmatrix}_{x=x^{*}(\mu)}))),$
where $eigvals$ refers to the operation of computing eigenvalue and $Re$ refers to the operation of taking the real part. Note that when the recovery rate of a system shifts from negative value to positive value, the bifurcation occurs(7). In other words, if we know the relation between the recovery rate of a system and the bifurcation parameter, we can easily identify the tipping point for this system. It is also worth noting that the relation between the recovery rate and the bifurcation parameter is the same for the systems of the same bifurcation type in normal forms, as shown in Fig. 2 D-F respectively. We assume that our DL algorithm can detect features of the recovery rate for the system by training it on a training set generated from a sufficiently diverse library of possible dynamical systems with fold, Hopf and transcritical bifurcation. Then the DL algorithm can be used to predict the tipping points where the recovery rate becomes zero based on the detected relation between the recovery rate and the bifurcation parameter of the system. Note that we only focus on dynamical systems exhibiting fold, Hopf and transcritical bifurcation in this paper. In order to predict tipping points of other bifurcation types, we can expand the training library by including simulated data exhibiting those dynamics.
Fig. 2. (A-C) The normal forms of fold, (supercritical) Hopf, and transcritical bifurcation where bifurcation occurs at $\mu=0$ (hollow square). Their mathematical expressions are inside the boxes within the corresponding figures. (D-F) The recovery rate $\lambda$ of the normal forms for the fold, (supercritical) Hopf, and transcritical bifurcation as a function of the control parameter $\mu$, their function expressions are inside the boxes within the corresponding figures. Bifurcation occurs when the recovery rate $\lambda$ reaches zero (hollow square).
Performance of DL algorithm on model time series
We applied the DL model on regularly-sampled time series generated from three ecological models with white noise and three climate models with red noise. We generated test data for each model by changing the bifurcation parameter with eleven different initial values and five different changing rates. We generated 10 test time series for each initial value and changing rate of the bifurcation parameter (50 test time series for each initial value). To evaluate the performance of DL algorithm, we measured the relative error(25) of tipping points prediction for each of the 50 test time series associated with every initial value of the bifurcation parameter. The performance of DL algorithm is evaluated by the mean relative error of tipping points prediction, which is averaged over the 50 measurements of the relative error of tipping points prediction for each initial value of the bifurcation parameter. We designed an ablation study by dropping out the 2D CNN layer, with LSTM serving as a competing algorithm. We first compared the results of DL algorithm with those of degenerate fingerprinting(9), DEV(12) and LSTM in three ecological models with white noise which exhibit fold, Hopf and transcritical bifurcations respectively. Then we compared the results of DL algorithm with those of BB method(10), DEV(12) and LSTM in three climate models with red noise which exhibit fold, Hopf and transcritical bifurcations respectively. As Fig. 3 shows, the DL algorithm outperforms the other competing algorithms for all initial values of each model. Moreover, the DL algorithm exhibits smaller fluctuations of relative error of tipping points prediction than the other competing algorithms for each initial value and exhibits smaller fluctuations of the mean relative error of tipping points prediction than the other competing algorithms across different initial values. Our results suggest that the DL algorithm is more robust against different initial values of the bifurcation parameter than competing algorithms.
We further tested the DL model on irregularly-sampled time series from the above six models. It is worth noting that the above competing algorithms are not applicable for irregularly-sampled time series. Therefore, we used linear interpolation to transform these irregularly-sampled time series into equidistant data. This allows us to use the competing algorithms of degenerate fingerprinting, BB method and DEV for detecting early warning signals on reconstructed model time series(42). LSTM is also applied as a competing algorithm. From Fig. 4, we find that DL algorithm outperforms the other competing algorithms for all initial values of each model. Moreover, our results suggest that the DL algorithm is more robust against different initial values of the bifurcation parameter than competing algorithms.
Fig. 3. We tested the DL model on regularly-sampled model time series with eleven initial values of bifurcation parameter. Here the mean relative error of tipping points prediction is plotted as line graphs against the initial value of the parameter. The area covered by the polyline represents the 90% confidence interval for the relative error of tipping points prediction. (a-c) We compared the DL algorithm (red lines) with degenerate fingerprinting (blue lines), DEV (green lines) and LSTM (purple lines) on three ecological model time series with white noise. These model time series undergo fold, Hopf, and transcritical bifurcation, respectively. (d-f) The DL algorithm (red lines) is compared with BB method (blue lines), DEV (green lines) and LSTM (purple lines) on three climate model time series with red noise. These model time series undergo fold, Hopf, and transcritical bifurcation, respectively.
In the ablation study, our DL algorithm achieves a mean relative error about 9% in predicting tipping points on these model time series, while LSTM exhibits a mean relative error about 23%, as shown in Supplementary information Fig. S1 and Fig. S2. These results suggest that the accuracy of predicting tipping points is higher when the LSTM layer receives sequences of features from the reconstructed system, rather than sequences of a single variable of the system. We also plotted the line graphs of the mean relative error of tipping points prediction against the distance between the final value of bifurcation parameter time series and the value of the tipping point, as shown in Supplementary information Fig. S3 and Fig. S4. Our results suggest that the DL algorithm is more robust against different distance to the tipping point than competing algorithms. Moreover, we plotted the line graphs of the mean relative error of tipping points prediction against the initial value of the parameter in five different changing rates of the bifurcation parameter, as shown in Supplementary information Fig. S5 and Fig. S6. Our results suggest that the DL algorithm is robust against various changing rates of bifurcation parameter.
Fig. 4. We tested the DL model on irregularly-sampled model time series with eleven initial values of bifurcation parameter. Here the mean relative error of tipping points prediction is plotted as line graphs against the initial value of the parameter. The area covered by the polyline represents 90% confidence interval for the relative error of tipping points prediction. (a-c) We compared the DL algorithm (red lines) with degenerate fingerprinting (blue lines), DEV (green lines) and LSTM (purple lines) on three ecological model time series with white noise. (d-f) The DL algorithm (red lines) is compared with BB method (blue lines), DEV (green lines) and LSTM (purple lines) on three climate model time series with red noise. We used linear interpolation to transform these irregularly-sampled time series into equidistant data so that they are suitable for degenerate fingerprinting, BB method and DEV.
We then tested the DL model on irregularly-sampled time series from the ascending arousal system with a fold/fold hysteresis loop. We conducted irregular sampling of 400 points from twenty-two initial values of the bifurcation parameter, which are $0.1,0.2,\dots,1.1$ (bifurcation parameter increases from these initial values) and $1.9,1.8,\dots,0.9$ (bifurcation parameter decreases from these initial values). We applied the DL model on these irregularly-sampled time series and the mean relative errors of tipping points prediction are 3.12% and 3.31% respectively, for the increasing and decreasing bifurcation parameter cases, as shown in Fig. 5 (a). Similarly, we tested the DL model on irregularly-sampled time series from the Sprott B system with a Hopf/Hopf-hysteresis bursting. We irregularly sampled 400 points from twenty-two initial values of the bifurcation parameter, which are $\pi,1.04\pi,\dots,1.4\pi$ (bifurcation parameter increases from these initial values) and $2\pi,1.96\pi,\dots,1.6\pi$ (bifurcation parameter decreases from these initial values). We applied the DL model on these irregularly-sampled time series and the mean relative errors of tipping points prediction are 2.82% and 2.54% respectively, for the increasing and decreasing bifurcation parameter cases, as shown in Fig. 5 (b). Our results suggest that the DL algorithm is effective for tipping point prediction in theoretical models with hysteresis phenomenon.
Fig. 5. The performance of the DL algorithm in predicting transitions of irregularly-sampled time series generated by theoretical models exhibiting hysteresis. The red and blue curves illustrate the time series forced by increasing and decreasing bifurcation parameter respectively. The orange dots denote the DL predictions, and the orange short vertical lines denote the initial points of the time series data used for prediction. We connect them with orange lines. (a) In the case of the sleep-wake system, a fold/fold hysteresis loop is observed at $D=1.153$ (red dashed line) and $D=0.883$ (blue dashed line) respectively. The predicted tipping points are between $[1.148,1.169]$ and $[0.8679,0.8898]$ respectively. (b) For the Sprott B system, a Hopf/Hopf hysteresis bursting occurs at $k=1.461\pi$ (red dashed line) and $k=1.539\pi$ (blue dashed line) respectively. The predicted tipping points are between $[1.454\pi,1.464\pi]$ and $[1.534\pi,1.545\pi]$ respectively.
Performance of DL algorithm on empirical time series
We tested the DL model on two empirical examples, including a cyanobacteria microcosm experiment under light stress(38) and a physical experiment of voice oscillation during heat(20). For each empirical system, we applied the DL model to several irregularly-sampled time series. Each time series includes 400 data points with different initial values of the control parameter. Then we used linear interpolation to transform these irregularly-sampled records to time series with equidistant data. This allows us to use the competing algorithms of degenerate fingerprinting, BB method and DEV for detecting early warning signals on reconstructed empirical records(42). Note that these two empirical examples from ecology and physics are subject to white noise. Therefore, we applied degenerate fingerprinting to these empirical data and did not use the BB method for comparison. We compared the results of the DL algorithm with two competing algorithms, which are degenerate fingerprinting and DEV.
In the cyanobacteria microcosm experiment under light stress, the photo-inhibition drives a cyanobacterial population, measured by light attenuation coefficient, towards a tipping point with fold bifurcation when a critical threshold of light irradiance is approached(38). We sampled seven time series with different initial values of light irradiance which are 477, 517, 557, 903, 944, 985 and 1025 $\rm\mu$mol photons $\rm m^{-2}s^{-1}$. Since the data between 879 and 903 $\rm\mu$mol photons $\rm m^{-2}s^{-1}$ is missing, the time series is sampled with initial values of 477, 517, 557 $\rm\mu$mol photons $\rm m^{-2}s^{-1}$ respectively and the same final value of 879 $\rm\mu$mol photons $\rm m^{-2}s^{-1}$. The mean relative error of the prediction with our DL algorithm is 3.82%, where the ground truth of the tipping point is 1091 $\rm\mu$mol photons $\rm m^{-2}s^{-1}$. As shown in Fig. 6 A1-A3, the DL algorithm outperforms the other competing algorithms in all cases. These results suggest the robustness of the DL algorithm in tipping points prediction against different initial values of light irradiance.
The second empirical data we analyzed is the thermoacoustic system. The state of the thermoacoustic system is measured by acoustic pressure. The thermoacoustic system undergoes a Hopf bifurcation from a non-oscillatory to an oscillatory state with increasing voltage(24, 43, 44, 45). It has also been demonstrated that the transition to high amplitude limit cycle oscillations occurs later for faster changing rate of voltage, which can be explained by rate-delayed tipping(20, 21). We sampled time series with nine initial values of voltage which are $0,0.2,\dots,1.6$ V under voltage changing rates of 20 mV/s and 40 mV/s, and with ten initial values of voltage which are $0,0.2,\dots,1.8$ V under a voltage changing rate of 60 mV/s. We then applied a change-point detection algorithm(46) to detect tipping points of these empirical time series under voltage changing rates of 20 mV/s, 40 mV/s and 60 mV/s, which are 1.72 V, 1.76 V and 1.87 V respectively. As shown in Fig. 6 B1-B3, we find that the DL algorithm outperforms the other competing algorithms in predicting the tipping points in most cases under the voltage changing rate of 40 mV/s, with mean relative error of 4.31% by our DL algorithm. We further test these algorithms for other two changing rates of voltage as shown in Supplementary information Fig. S7, which suggests the DL algorithms also shows best performance among all these algorithms. These results suggest that the DL algorithm is robust against various initial values and changing rages of voltage.
Discussion
Predicting the occurrence of tipping points based on time series is a challenging problem. In this work, we develop a deep learning algorithm that exploits information of normal form from the training systems for tipping points prediction. This algorithm can deal with both regularly-sampled and irregularly-sampled time series by using the time series sampled from a single variable of a system. Our results show that the DL algorithm not only outperforms traditional methods in the accuracy of tipping points prediction for regularly-sampled time series, but also achieve accurate prediction for irregularly-sampled time series. Our results pave the way to make effective strategies to prevent and prepare tipping points from various real-world systems(47).
There are three major advantages of our work. First, traditional methods for tipping points prediction can only deal with regularly-sampled time series(9, 10, 11, 12). However, the DL algorithm used in this paper can deal with both regularly-sampled time series and irregularly-sampled time series. This is because the convolution kernels in the 2D CNN layer can extract features of the reconstructed system from segments of irregularly-sampled time series (Supplementary information Note S3), which is used for tipping points prediction in the LSTM layer. Second, most existing methods(9, 48, 25, 26, 27) require information of multiple interacting variables in the study system for predicting tipping points. However, based on the embedding theorem for irregular sampling(28), the DL algorithm only requires the time series sampled from a single variable of the study system. Third, the effect of rate-delayed tipping has been ignored in previous methods for tipping points prediction(9, 10, 11, 12). This is due to the limitations of the theory of fast-slow systems used in previous methods. However, we consider the effect of rate-delayed tipping in our work and label the data by using the fact that the quasi-static attractor loses stability at the occurrence of tipping points in our training set (see Methods for details). The DL algorithm trained with such labels is robust against different changing rates of bifurcation parameter (Supplementary information Figs. S5-S7).
The embedding theorem for irregular sampling(28) requires that the dimension $d$ of the system reconstructed by delay embedding must be larger than twice the dimension $m$ of the study system (Supplementary information Note S3). Therefore, the length of convolution kernels used in the CNN layer of our DL algorithm is required to be larger than $2m$. However, due to the time-varying nonstationarity of dynamical system approaching a bifurcation, the features extracted from shorter convolutional kernel should contain much more dynamical information of the system, compared to those extracted from longer convolutional kernel(49). Therefore, our DL algorithm may not perform well in predicting tipping points in high-dimensional systems. Another limitation of our DL algorithm is that it requires the existence of tipping points as prior knowledge. If there is no tipping point in the system, the output of the DL algorithm is meaningless. Therefore, several methods such as DEV(12) or some DL classifiers(18, 22, 23, 24) should be used to determine whether a system is approaching a tipping point first. Then our DL algorithm can be applied to predict the specific location of a tipping point if it exists.
Fig. 6. The performance of DL algorithm in predicting tipping points on irregularly-sampled empirical time series. The dots denote the DL predictions, and the short vertical lines denote the initial points of the time series data used for prediction. We connect them with lines. (A1-A3) Cyanobacterial population undergoes a fold bifurcation, which is measured by light attenuation coefficient. The fold bifurcation occurs at 1091 $\rm\mu$mol photons $\rm m^{-2}s^{-1}$, while the predictions by our DL algorithm are between 1074 $\rm\mu$mol photons $\rm m^{-2}s^{-1}$ and 1106 $\rm\mu$mol photons $\rm m^{-2}s^{-1}$. (B1-B3) The thermoacoustic system undergoes a Hopf bifurcation under the 40 mV/s changing rate of the control parameter voltage. The Hopf bifurcation occurs at 1.76 V, while the predictions by our DL algorithm are between 1.71 V and 1.79 V. We compared the performance of the DL algorithm (red) with degenerate fingerprinting (blue) and DEV (green). We used linear interpolation to transform these irregularly-sampled time series into equidistant data so that they are suitable for degenerate fingerprinting and DEV.
To analyze whether the DL algorithm uses the recovery rate in the normal form for prediction, we designed two control experiments. First, we trained three DL models on three datasets, each consisting solely of time series with fold, Hopf, and transcritical bifurcation, respectively. Then we applied these three DL models on irregularly-sampled model time series from six theoretical models. We found that the DL model trained on dataset containing the features of a bifurcation perform better in predicting tipping points of test time series with that bifurcation, compared to another two trained on datasets without the corresponding normal form features (Supplementary information Fig. S8). These results suggest that the DL algorithm can extract the features of normal form. Second, we trained two DL models on two datasets, each consisting solely of time series with supercritical and subcritical pitchfork bifurcation, respectively (Supplementary information Note S7). Then we applied these two DL models on irregularly-sampled model time series from a model with supercritical pitchfork, where both models achieved similar performance in predicting the tipping points (Supplementary information Fig. S9). The normal forms of supercritical and subcritical pitchfork bifurcations exhibit the same relation between the recovery rate and the bifurcation parameter, and only differ in the cubic term. Thus, the DL model trained on time series with subcritical pitchfork bifurcation can be used to predict tipping points of time series with supercritical pitchfork bifurcation, which suggest that the DL algorithm can extract the features of the recovery rate in normal form.
Our work raises several problems worthy of future pursuit. First, we have studied the situation of local codimension-one bifurcations in continuous-time dynamical systems in this paper. For future work, we can focus on tipping points prediction in discrete-time dynamical systems, such as period-doubling bifurcation(7) which arise naturally in physiology(50, 51) and ecology(52). It would be interesting to develop a method that can deal with both regularly-sampled and irregularly-sampled time series for systems with period-doubling bifurcation. Second, it would be interesting to investigate tipping points prediction in systems with other types of bifurcation, such as codimension-two bifurcation and global bifurcation(7). Yet, tipping points prediction in systems with codimension-two bifurcation and global bifurcation are more challenging than that in systems with codimension-one bifurcation. Third, one can develop interpretable machine learning models for tipping points prediction by combing dynamical system theory with neural networks(53, 54, 55, 56, 57), which offer an avenue for making safe and reliable high-stakes decisions for policy makers(58).
Methods
Generation of training data for the DL algorithm
We construct two-dimensional dynamical systems of the following form which are used to generate the training data based on simulation(18)
$\frac{dx}{dt}=\sum\limits_{i=1}^{10}a_{i}p_{i}(x,y)$ (1)
$\frac{dy}{dt}=\sum\limits_{i=1}^{10}b_{i}p_{i}(x,y),$
where $x$ and $y$ are state variables, $p(x,y)$ is a vector containing all polynomials in $x$ and $y$ from zero up to third order
$p(x,y)=(1,x,y,x^{2},xy,y^{2},x^{3},x^{2}y,xy^{2},y^{3}),$
$p_{i}(x,y)$ is the $i$-th component of $p(x,y)$. $a_{i}$ and $b_{i}$ are parameters randomly drawn from standard normal distribution, and then, half of these parameters are selected at random and set to zero. The parameters for the cubic terms are set to the negative of their absolute values to encourage models with bounded solutions. Since our training data is required to contain representation of all possible dynamics that may occur in the study time series data, we generate models with different sets of parameter values in Eqs. 1 until a required number of each type of bifurcation (fold, Hopf, or transcritical) has been discovered. We add white or red noise to these models to simulate training data for the DL algorithm.
After a model is generated, we perform numerical simulation of this model with 10,000 time steps from a randomly drawn initial condition and test whether the system converges to an equilibrium point. The odeint function from the Python package Scipy(59) is applied in the numerical simulation with a step size of 0.01. The criteria that we used to determine the convergence is that the maximum difference between the final 10 points in numerical simulation is less than $10^{-8}$. The convergence is required in order to search for bifurcations which occur at non-hyperbolic equilibria(7). The models that do not converge are discarded. For the models that converge, we apply AUTO-07P(60) to identify bifurcations along the equilibrium branch by either increasing or decreasing each nonzero parameter. For each bifurcation identified, we first set the initial condition with the value of the equilibrium of the model and a burn-in period of 100 units of time. Then we run simulations of the model with noise and obtain the quasi-static attractor time series with noise and bifurcation parameter time series which are used for training. We also run simulations of the same model without noise and obtain the corresponding quasi-static attractor time series and bifurcation parameter time series which are used for calculating the recovery rate $\lambda$ to locate the bifurcation point where $\lambda$ changes from negative value to positive value,
$\lambda=max(Re(eigvals(\begin{bmatrix}\frac{\partial(\sum\limits_{i=1}^{10}a_{i}p_{i}(x,y))}{\partial x}&\frac{\partial(\sum\limits_{i=1}^{10}a_{i}p_{i}(x,y))}{\partial y}\\ \frac{\partial(\sum\limits_{i=1}^{10}b_{i}p_{i}(x,y))}{\partial x}&\frac{\partial(\sum\limits_{i=1}^{10}b_{i}p_{i}(x,y))}{\partial y}\end{bmatrix}_{(x,y)=(x^{*},y^{*})}))).$ (2)
The simulations of the models with noise are based on the Euler Maruyama method(61) with a step size of 0.01 and the simulations of the models without noise are based on Euler method(62) with the same step size.
For each model where a bifurcation is identified, we perform five independent simulations of the model with varying bifurcation parameter from its initial value to a terminal value beyond the bifurcation point given by AUTO-07P. In these five simulations, we set the changing rate of the bifurcation parameter to remain constant within each simulation where the changing rate of the bifurcation parameter is the change in the bifurcation parameter per unit time with a step size of 0.01, but vary across different simulations. The ratio of each changing rate of the bifurcation parameter to the smallest changing rate of the bifurcation parameter in these five simulations is drawn from $1,2,\dots,10$. The nonzero changing rate of the bifurcation parameter can lead to the delay in the occurrence of tipping points, which is called rate-delayed tipping(20, 21). Due to rate-delayed tipping, the theoretical bifurcation point given by AUTO-07P is not the accurate location of the tipping point. Therefore, we identify the location of the tipping point by the recovery rate (Eq. 2) and use this identified tipping point as the training label for our supervised deep learning training process (Supplementary information Note S2). If no tipping point is detected in a simulation, this simulated data will be discarded. Otherwise, we randomly and irregularly sample $l_{s}$ points from quasi-static attractor time series with noise which is simulated by the variable $x$ of the model and corresponding bifurcation parameter time series between the initial value and the tipping point, where $l_{s}$ is drawn from $Uniform(505,1000)$. Then we select the first 500 points as training data. The obtained training data can make DL algorithm to learn the features of recovery rate from time series data with different changing rates of parameter and varying distances between the end of the time series and the tipping points.
Here we train deep learning algorithm on datasets with white and red noise. The white noise in simulation has the amplitude drawn from a triangular distribution centered at 0.01 with upper and lower bounds 0.0125 and 0.0075 respectively while the red noise is modeled by AR(1) process in which the lag-1 autoregressive coefficient is between -1 and 1. It is worth noting that the simulations shift to new regime before the bifurcation point due to the noise(63, 64). Since the location of this premature transition is stochastical, we still use the location where the quasi-static attractor without noise loses stability as the training label for DL algorithm even it is not accurate. We mathematically illustrate how slight perturbation causes bifurcation to occur earlier and its stochasticity in Supplementary information Note S6.
DL algorithm architecture and training
In this paper we use the 2D CNN-LSTM DL algorithm(65, 66). The 2D CNN-LSTM architecture is shown in Fig. 1. We use a train/validation/test split of 0.95/0.04/0.01 for the DL algorithm, and MSE as the loss function to minimize the difference between the real tipping points (training labels) and the predictions. The DL algorithm is trained on 300,000 instances for 200 epoches. We train ten neural networks and report the performance of the DL algorithm averaged over them. The hyperparameters of the DL model are tuned through random search, and we present the optimal hyperparameters of the DL model in Supplementary information Table S1.
Before training our DL algorithm, there are several preprocessing matters(67). The quasi-static attractor time series with noise are detrended using Lowess smoothing with a span of 0.2 to obtain the residual time series used for training(68). Besides, in order to obtain the robustness of DL algorithm on time series of various lengths, we zero out the left side of the residual time series and corresponding bifurcation parameter time series, with the length drawn from $Uniform(0,250)$. Due to the zeroing process, the DL algorithm can deal with time series even if it is a little short. After the zeroing process, each residual time series is normalized by dividing each time series data point by the average absolute value of the residuals across the non-zero part of the time series. In addition, each bifurcation parameter time series is also normalized: each data point in the time series minus the initial value of the bifurcation parameter time series, and then divided by the distance between the initial and final values of the bifurcation parameter time series, thereby mapping the bifurcation parameter time series to the interval of $[0,1]$. The corresponding label, i.e., the tipping point, is normalized following the same procedure as the normalization of the data point in bifurcation parameter time series. Thus the relative distance of tipping point to the time series in the training set for the DL algorithm (calculated as the tipping points minus the final value of the bifurcation parameter time series, divided by the distance between the initial and final values of the bifurcation parameter time series) is between 0.01 and 2. In other words, the labels in the training set, after normalization, will fall within the interval of $[1.01,3]$.
Theoretical models used for testing
The simulation of theoretical models with noise is based on the Euler Maruyama method(61) with a step size of 0.01 ($\delta t=0.01$ units of time) unless otherwise stated. The amplitude of white noise $\sigma$ is 0.01 while the red noise is modeled by AR(1) process in which the lag-1 autoregressive coefficient $\phi$ is drawn from $Uniform(-1,1)$. We set five different changing rates of the bifurcation parameter in simulations of each theoretical model. The bifurcation point of each simulation is the location where the recovery rate changes from negative value to positive value. Then we sample from these simulations to obtain regularly-sampled model time series and irregularly-sampled model time series for testing. The lengths of these model time series range from 250 to 500, and the relative distances of tipping points to these time series range from 0.01 to 2.
We apply the relative error $\varepsilon$ of tipping points prediction to evaluate the performance of the DL algorithm, which is defined as
$\varepsilon=\frac{\left|\widehat{\mu}_{c}-\mu_{c}\right|}{\left|\mu_{end}-\mu_{c}\right|},$
where $\widehat{\mu}_{c}$ is the predicted tipping point, $\mu_{c}$ is the real tipping point, $\mu_{end}$ is the value of bifurcation parameter for last point of the test time series.
Theoretical models with white noise
To test the fold bifurcation with white noise, we use the May’s harvesting model(29). This is given by
$\frac{dx}{dt}=rx(1-\frac{x}{k})-\frac{hx^{2}}{s^{2}+x^{2}}+\sigma\xi(t),$
where $x$ represents biomass of some population, $k$ is its carrying capacity, $h$ is the harvesting rate, $s$ characterizes the nonlinear relationship between harvesting output and current biomass, $r$ is the intrinsic per capita growth rate of the population, and $\xi(t)$ is a Gaussian white noise process. We use parameter values $r=1$, $k=1$, $s=0.1$ and $h$ increases at rates of $1\times 10^{-5},2\times 10^{-5},\dots,5\times 10^{-5}$. We generate model time series by this equation from eleven initial values of $h$, which are $0,0.02,0.04,\dots,0.18,0.2$. In this configuration, the fold bifurcation occurs in $h\in[0.2713,0.2926]$.
To test the Hopf bifurcation with white noise, we use a three-species chaotic food chain model(30). This is given by
$\frac{dr}{dt}=r(1-\frac{r}{k})-\frac{x_{c}y_{c}cr}{r+r_{0}}+\sigma\xi_{r}(t),$
$\frac{dc}{dt}=x_{c}c(\frac{y_{c}r}{r+r_{0}}-1)-\frac{x_{p}y_{p}pc}{c+c_{0}}+\sigma\xi_{c}(t),$
$\frac{dp}{dt}=x_{p}p(\frac{y_{p}c}{c+c_{0}}-1)+\sigma\xi_{p}(t),$
where $r$, $c$, and $p$ represent the population densities of the resource, consumer, and predator species, respectively. $k$ characterizes the environmental capacity of the resource species. $x_{c}$, $y_{c}$, $x_{p}$, $y_{p}$, $r_{0}$, and $c_{0}$ are other parameters in the system. $\xi_{r}(t)$, $\xi_{c}(t)$ and $\xi_{p}(t)$ are independent Gaussian white noise processes. For our simulations, we use parameter values $x_{c}=0.4$, $y_{c}=2.009$, $x_{p}=0.08$, $y_{p}=2.876$, $r_{0}=0.16129$, $c_{0}=0.5$ and $k$ increases at rates of $1\times 10^{-5},2\times 10^{-5},\dots,5\times 10^{-5}$. We generate model time series by these equations from eleven initial values of $k$, which are $0.20,0.22,0.24,\dots,0.38,0.40$. In this configuration, the Hopf bifurcation occurs in $k\in[0.4656,0.4839]$.
To test the transcritical bifurcation with white noise, we use the Rozenzweig-MacArthur consumer-resource model(31). This is given by
$\frac{dx}{dt}=gx(1-\frac{x}{k})-\frac{axy}{1+ahx}+\sigma\xi_{x}(t),$
$\frac{dy}{dt}=\frac{eaxy}{1+ahx}-my+\sigma\xi_{y}(t),$
where $x$ and $y$ represent the population densities of the resource, consumer, respectively. $g$ is the intrinsic per capita growth rate of $x$, $k$ is its carrying capacity, $a$ is the attack rate of $y$, $e$ is the conversion factor, $h$ is the handling time, $m$ is the per capita consumer mortality rate, $\xi_{x}(t)$ and $\xi_{y}(t)$ are independent Gaussian white noise processes. We fix the parameter values $r=4$, $k=1.7$, $e=0.5$, $h=0.15$, $m=2$ and $a$ increases at rates of $1\times 10^{-4},2\times 10^{-4},\dots,5\times 10^{-4}$. We generate model time series by these equations from eleven initial values of $a$, which are $0,0.5,1.0,\dots,4.5,5.0$. In this configuration, the transcritical bifurcation in $a\in[5.8819,5.8823]$.
Theoretical models with red noise
To test the fold bifurcation with red noise, we use a climate model describing temperature of an ocean on a spherical planet subjected to radiative heating(33, 32) which can simulate a transition from a greenhouse to an icehouse Earth(42). The model is given by
$\frac{dT}{dt}=\frac{-e\rho T^{4}+\frac{1}{4}uI_{0}(1-a_{p})}{c}+\eta(t),$
$\text{with}\quad a_{p}=a-bT,$
where $T$ represents the average temperature of ocean, $u$ is relative intensity of solar radiation, $e$ is effective emissivity, $I_{0}$ is solar irradiance, $c$ is a constant thermal inertia, and $a_{p}$ is the planetary albedo. $a$ and $b$ are parameters which define a linear feedback between ice and albedo variability and temperature. $\eta(t)$ is a red noise process with lag-1 autoregressive coefficient $\phi$. We use $e=0.69$, $I_{0}=71944000$, $c=10^{8}$, $a=2.8$, $b=0.009$, $\rho=0.03$, $\delta t=1$ time unit and $u$ decreases at rates of $5\times 10^{-7},6\times 10^{-7},\dots,9\times 10^{-7}$. We generate model time series by this equation from eleven initial values of $u$, which are $1.4,1.38,1.36,\dots,1.22,1.2$. In this configuration, the fold bifurcation occurs in $u\in[0.9596,0.9631]$.
To test the Hopf bifurcation with red noise, we use the middle Pleistocene transition system(34), explaining the dynamics of glacial cycles, which is given by
$\frac{dx}{dt}=-x-y+\eta_{x}(t),$
$\frac{dy}{dt}=-pz+uy+sz^{2}-yz^{2}+\eta_{y}(t),$
$\frac{dz}{dt}=-q(x+z)+\eta_{z}(t),$
where $x$, $y$ and $z$ represent the global ice volume, the atmospheric greenhouse gas concentration and the deep ocean temperature, respectively. All variables are rescaled to dimensionless form, $p=1$, $q=1.2$, $s=0.8$ are parameters, and $u$ increases at rates of $1\times 10^{-5},2\times 10^{-5},\dots,5\times 10^{-5}$. $\eta_{x}(t)$, $\eta_{y}(t)$ and $\eta_{z}(t)$ are independent red noise processes with lag-1 autoregressive coefficient $\phi$. We generate model time series by these equations from eleven initial values of $u$, which are $0,0.03,0.06,\dots,0.27,0.3$. In this configuration, the Hopf bifurcation occurs in $u\in[0.3546,0.3547]$.
To test the transcritical bifurcation with red noise, we use the simplified version of TRIFFID dynamic global vegetation model(35, 69). It can be used to simulate the dieback of the Amazon rainforest(69), which is mainly driven by climate change(70). The model is given by
$\frac{dV}{dt}=PV^{*}(1-V)-GV+\eta(t),$
where $V$ represents the broadleaf fraction, $G$ is a disturbance coefficient (0.004/year) and $V^{*}$ is either the value of $V$ or 0.1 if $V$ falls below 0.1. $\eta(t)$ is a red noise process with lag-1 autoregressive coefficient $\phi$. $P$ is the productivity, in dimensionless area fraction units, it decreases at rates of $1\times 10^{-5},2\times 10^{-5},\dots,5\times 10^{-5}$. We generate model time series by this equation from eleven initial values of $P$, which are $0.90,0.82,0.74,\dots,0.18,0.10$. In this configuration, the transcritical bifurcation occurs in $P\in[-0.0061,-0.0046]$.
Theoretical models with hysteresis phenomenon
We also test our DL model on irregularly-sampled model time series generated by two theoretical neuroscience models with white noise. These systems exhibit hysteresis, which suggests that when the bifurcation parameter changes in opposite directions, these system will undergoes sudden shifts at different bifurcation points. The ascending arousal system is modeled in terms of the neuronal populations and their interactions(36), which is given by
$\tau_{v}\frac{dV_{v}}{dt}=-V_{v}+v_{vm}Q_{m}+D+\sigma\xi_{v}(t),$
$\tau_{m}\frac{dV_{m}}{dt}=-V_{m}+v_{ma}Q_{a}+v_{mv}Q_{v}+\sigma\xi_{m}(t).$
[…]
Another system is Sprott B system(71) with a single excitation, which can be used for researching bursting oscillation in neuroscience(37). The model is given by
$\frac{dx}{dt}=a(y-x)+\sigma\xi_{x}(t),$
$\frac{dy}{dt}=xz+\beta\cos(k)+\sigma\xi_{y}(t),$
$\frac{dz}{dt}=b-xy+\sigma\xi_{z}(t),$
where $a=8$, $b=2.89$ and $\beta=5$. $\xi_{x}(t)$, $\xi_{y}(t)$ and $\xi_{z}(t)$ are independent Gaussian white noise processes. We force the parameter $k$ to increase from $\pi$ or decrease from $2\pi$ at a rate of $\pi\times 10^{-3}$ within the interval of $[\pi,2\pi]$. A Hopf/Hopf-hysteresis bursting occurs at $k=1.461\pi$ (in the increasing direction) and $k=1.539\pi$ (in the decreasing direction) respectively in this configuration.
Empirical systems used for testing
In this work, we use two empirical examples in the fields of microbiology and thermoacoustics to evaluate the performance of the DL algorithm.
In the first example, photo-inhibition drives cyanobacterial population to a fold bifurcation when a critical light level is exceeded(38). In this system, bifurcation parameter light irradiance starts at 477 $\rm\mu$mol photons $\rm m^{-2}s^{-1}$ and is increased in steps of 23 $\rm\mu$mol photons $\rm m^{-2}s^{-1}$ each day. The time series includes 7,784 data points spanning overall 28.86 days with time interval equal to 0.0035 day (5 min).
The second example is a thermoacoustic system where positive feedback between the unsteady heat release rate fluctuations and the acoustic field in the confinement can result in a transition to high amplitude limit cycle oscillations(45). Induja et al.(20) conduct experiments in a thermoacoustic system exhibiting Hopf bifurcation. They pass several constant mass flow rates of air through a horizontal Rijke tube which consists of an electrically heated wire mesh in a rectangular duct and increase the voltage applied across the wire mesh to attain the transition to thermoacoustic instability. In such case, they observe that the occurrence of Hopf bifurcation depends on the changing rate of control parameter. We select three experimental sets for our study with the bifurcation parameter of voltage increasing at rates of 20 mV/s, 40 mV/s, and 60 mV/s. The voltage is ranging from 0 V to 2.4 V. The corresponding time series data consists of 1,200,000, 600,000, and 400,000 data points, respectively.
Comparison of DL algorithm with competing algorithms
For detecting early warning signals (EWS) in quasi-static attractor time series with white noise, H. Held and T. Kleinen(9) develop a technique called degenerate fingerprinting using PCA to approximate the critical mode, then estimating the lag-1 autoregressive coefficient of critical mode as an indicator (Supplementary information Note S5). Boettner and Boers(72, 10) propose an unbiased estimate of the lag-1 autoregressive coefficient for time series with red noise which we refer to as the BB method (Supplementary information Note S5). Florian Grziwotz et al.(12) introduce a EWS robust to limited level of noise, named dynamical eigenvalue (DEV), which is rooted in bifurcation theory of dynamical systems to estimate the dominant eigenvalue of the system (Supplementary information Note S5). After obtaining EWS through these three approaches, linear regression or nonlinear fit between EWS and bifurcation parameters is used to make an extrapolation to anticipate the occurrence of tipping points. In our comparative experiment, we employ linear regression or fit a quadratic curve for extrapolation, and then we select the estimate with the better performance.
It is worth noting that the required information of detecting early warning signals with degenerate fingerprinting is different from that with our DL algorithm. The degenerate fingerprinting requires information from all variables of the study system to approximate the critical mode, whereas our DL algorithm only requires data from one variable of the study system. Similarly, the DEV also requires data from one variable of the study system. Moreover, the BB method is designed for estimating the lag-1 autoregressive coefficient of one-dimensional system, whereas our DL algorithm is suitable for analyzing multidimensional system. The DEV can also deal with multidimensional system.
Data availability
All the simulated datasets and the empirical datasets used to test the deep learning algorithm have been deposited in Github (https://github.com/zhugchzo/dl_occurrence_tipping) and the training set used to train the deep learning algorithm have been deposited in Zenodo (https://zenodo.org/records/13894933).
Code availability
Code and workflow to reproduce the analysis are available at the GitHub (https://github.com/zhugchzo/dl_occurrence_tipping).
Acknowledgments
This work is supported by National Natural Science Foundation of China under Grant 62388101, the National Key Research and Development Program of China under Grant 2022YFF0902800. We thank Didier Sornette, Jianxi Gao for valuable comments on this work.
Author contributions
W.C. conceived the project. All authors designed the project. C.Z. analyzed the empirical data and did the analytical and numerical calculations. All authors analyzed the results. C.Z. and W.C. wrote the manuscript.
Competing interests
The authors declare no competing interests.